HUBBLE SPACE TELESCOPE DAILY REPORT #1716 PERIOD COVERED: 0000Z (UTC) 09/11/96 - 0000Z (UTC) 09/12/96 Daily Status Report as of 256/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FGS 6874 (Parallaxes of Hyades Cluster Members, Cycle 6 = - High) The FGSs were used to make parallax astrometric observations of = the stars VA622 and VA627 and several other reference stars within the = Hyades star cluster, to better define the Population 1 zero age main = sequence. The observations were completed as planned, and no problems = were reported. 1.2 Completed Four Sets of WF/PC-2 6253 (Pure Parallel Archive = Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The observations were completed as planned, and no problems were = reported. 1.3 Completed FOC/96 6894 (Cycle 6: Aperture Monitor Location) The Faint Object Camera (f/96) was used to make observations of = selected targets in NGC-188 in order to monitor the location of the = camera's aperture with respect to the V2-V3 coordinate frame. The = observations were completed as planned, and no problems were reported. 1.4 Completed GHRS/2 6534 (Be-Deficient Halo Stars: Implications For = Galactic Chemical Evolution) The High Resolution Spectrograph (Side Two) was used to make = observations of the beryllium deficient halo star HD94028 to help study = the galactic chemical evolution. The objective is to measure the = abundance of boron in this star, and compare this with the abundances of = Be and Li. The observations were completed as planned, and no problems = were reported. 1.5 Completed FOS/RD 6292 (Time Resolve FOS Observations of the AP = Star HD60435) The Faint Object Spectrograph (Red Side) was used to make = observations of the star HD60435. During the SV program with the HSP = this star was found to have previously unknown pulsations. The = observations were completed as scheduled, and no anomalies were reported. 1.6 Completed FOS/RD 6796 (Spectral Mapping of Accretion Disks in = Cataclysmic Binaries) The Faint Object Spectrograph (Red Side) was used to make rapid = mode spectrographic mapping observations of the accretion disk around the = cataclysmic binary system V348-Puppis taken during the system's eclipse = phase. The observations were completed as planned, and no problems were = reported. 1.7 Completed GHRS/2 6644 (Boron Isotope Ratios in Early Stars from B = III) The High Resolution Spectrograph (Side 2) was used to make high = resolution spectrographic observations of early-type, B class star HD3346 = in order to determine its B11/B10 isotope ratio. Knowledge of the B = ratio will allow new constraints to be placed on theories of B11 = synthesis. The observations were completed as planned, and no problems = were reported. 1.8 Completed WF/PC-2 6363 (The Environments and Host Galaxies of = BL-Lac Objects) The WF/PC-2 was used to make snapshot survey observations of = BL-Lac objects MS1258.4+6 and MS0737.9+7 in order to study their = environment and their host galaxies. The goal is to investigate = systematic trends in host galaxy type, unusual morphological features, = nearby companion galaxies, etc. The observations were executed as = scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 5 Successful : 5 2.2 FHST Updates: Scheduled : 12 Successful : 12 2.3 Operations Notes: There were no reported operational problems. HST operations transferred to the SAMS backup facility in = preparation for today's simulation activities at approximately 256/0200Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: GSFC Internal Simulation #10 (EVA 2) is scheduled today from = approximately 8:00 a.m. until approximately 6:00 p.m. /DMH